WLCG Tier1 Service Coordination Minutes - 15 March 2012

Attendance

Local: Ian, Eva, Ale, Maite, Nicolo, Stefan, Julia, Jamie, MariaG, Mike Lamont, AndreaS, Claudio Grandi, Oliver, Maarten, MariaD, Daniel, Massimo, David.

Remote: FNAL, NDGF, BNL, RAL, PIC, IN2P3-CC, OSG, CNAF

Action list review

Release update

Data Management & Other Tier1 Service Issues

Site Status Recent changes Planned changes
CERN CASTOR (tapegateway active) + SRM-2.11 for all LHC instances; CASTOR 2.1.12-4 for ATLAS, CMS, LHCb and PUBLIC instances (w/ xroot-xcastor2fs_2112-1.1.0-1). CASTORALICE sports 2.1.11-9
FTS: all nodes in SLC5 3.7.0-3
EOSALICE, EOSATLAS and EOSCMS 0.1.2-2 (w/ xrootd-3.1)

EOSATLAS being rolled out

CASTORCMS+ATLAS+LHCb upgraded to 2.1.12-4

CASTORALICE to 2.1.12-4 (Mar 19th)
ASGC CASTOR 2.1.11-6
SRM 2.11-0
DPM 1.8.2-3
None None
BNL dCache 1.9.12.10 (Chimera, Postgres 9 w/ hot backup)
http (aria2c) and xrootd/Scalla on each pool
None None
CNAF StoRM 1.8.1 (Atlas, CMS, LHCb) None None
FNAL dCache 1.9.5-23 (PNFS, postgres 8 with backup, distributed SRM) httpd=2.2.3
Scalla xrootd 2.9.7/3.1.0.osg
Oracle Lustre 1.8.6
EOS 0.1.1-12/xrootd 3.1.0.osg with Bestman 2.0.10
FTS 3.7.7 on SL5
upgraded FTS yesterday  
IN2P3 dCache 1.9.12-16 (Chimera) on core servers and pool nodes.
New hardware (more RAM, SSD disks) for Chimera and SRM servers (with SL6).
Postgres 9.1
None None
KIT dCache
atlassrm-fzk.gridka.de: 1.9.12-11 (Chimera)
cmssrm-fzk.gridka.de: head nodes 1.9.5-26 (Chimera), pool nodes 1.9.5-6 through -25
gridka-dcache.fzk.de: head nodes 1.9.5-26 (PNFS), pool nodes 1.9.5-24,-25
xrootd (version 20100510-1509_dbg)
   
NDGF dCache 2.1 (Chimera) on core servers. Mix of 1.9.13 and 2.0.1 on pool nodes.    
NL-T1 dCache 1.9.12-10 (Chimera) (SARA), DPM 1.7.3 (NIKHEF)    
PIC dCache 1.9.12-14; PNFS on Postgres 9.0 None None
RAL CASTOR 2.1.11-8
2.1.11-8 (tape servers)
SRM 2.11-1
Just upgraded FTS None
TRIUMF dCache 1.9.5-28 with Chimera namespace None None

Other site news

CASTOR news

CERN operations and development

EOS news

xrootd news

dCache news

StoRM news

FTS news

  • FTS 2.2.8 EMI now running in production at: CERN on all the FTS servers (pilot, T2Export and T0Export), NDGF-T1, RAL-LCG2, Taiwan-LCG2, FZK-LCG2 (just yesterday, waiting confirmation everything is ok). Rollout to the T1s is ongoing.
  • Summary of 2.2.8 rollout - https://svnweb.cern.ch/trac/glitefts/wiki/FTS228RolloutPlanning
    • Detailed FTS server deployment plan: https://docs.google.com/spreadsheet/ccc?key=0AthhzXLQok7XdFpUeDBfLXE2S1RDZE4zcHp6QWVpUFE .
    • Alessandro from CNAF requested to postpone their date for FTS upgrade towards the end of the required period 28-29.3. Ale confirmed that 4 hrs max. are enough.
    • CMS (Nicolo) said 3 GGUS tickets were opened with FTS 2.2.8 experience recently. The new functionality of 'resume' transfers requires a patch, now ready and being tested on the pilot system in order to use a less frequent check-point (now checking every second and overloading the system). The new EMI release in April will contain this patch.
  • FTS 3 update by Oliver (slides on the agenda). Nxt demo on 21st March.

DPM news

LFC news

LFC deployment

Site Version OS, n-bit Backend Upgrade plans
ASGC NA NA NA NA
BNL 1.8.0-1 SL5, 64-bit Oracle None
CERN 1.8.2-0 64-bit SLC5 Oracle all servers are SLC5 64-bit virtual machines
CNAF 1.8.0-1 SL5 64-bit Oracle None
FNAL N/A     Not deployed at Fermilab
IN2P3 1.8.2-2 SL5 64-bit Oracle 11g  
KIT 1.7.4-7 SL5 64-bit Oracle Oracle backend migration pending
NDGF 1.7.4.7-1 Ubuntu 10.04 64-bit MySQL None
NL-T1 1.7.4-7 CentOS5 64-bit Oracle  
PIC 1.7.4-7 SL5 64-bit Oracle  
RAL 1.7.4-7 SL5 64-bit Oracle  
TRIUMF 1.7.3-1 SL5 64-bit MySQL None

Experiment issues

WLCG Baseline Versions

Status of open GGUS tickets

Review of recent / open SIRs and other open service issues

Conditions data access and related services

Database services

  • Experiment reports:
    • ALICE: Testing the use of Active Data Guard in order to replace Streams PVSS replication.
    • ATLAS: new schema ATLAS_CONF_TRIGGER_MC added to replication from CERN to Tier1 sites. ATLAS LFC replication to BNL being investigated.
    • CMS: Move to Active Data Guard (to replace Streams replication) is progressing. Streams apply handler errors due to changes in Oracle 11g.
    • Active Data Guard monitoring being improved.

  • Site reports:
Site Status, recent changes, incidents, ... Planned interventions
BNL ntr None
CNAF    
KIT Migrated the Compass-LFC schema to the new hardware (FTS-cluster) and upgraded this cluster to 11.2.0.3. The migration and upgrade process to 11g is finished at KIT/Gridka. Some problems with communication between FTS-frontend and FTS-DB observed. Probably related to SCAN-IP or SERVICE_NAME request. Working on it. None
IN2P3 Follow-up on upgrade issues with Oracle. Upgrade to 11g eventually next week - being agreed with experiments
PIC ntr None
RAL Follow-up on scan listener issues with Oracle, bug doc 1379498.1. Data Guard configuration running for CASTOR databases. None
SARA    
TRIUMF Compatible parameter updated to 11.2.0.3 None

AOB

-- AndreaSciaba - 14-Mar-2012

Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2012-03-15 - MariaDimou
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback